The force in the middle of two identical charges separated by 1.77 cm is equal to 70N. What is the magnitude of the two forces? 


1
Expert's answer
2022-03-25T09:11:02-0400

The Coulomb's law says


"F=k\frac{|q_1||q_2|}{r^2}"

The magnitude of charges


"q=\sqrt{\frac{Fr^2}{k}}=\sqrt{\frac{70*0.0177}{9*10^9}}=1.17*10^{-5}\:\rm C"